Find the characteristic root of the recurrence relation an=anà ƒ ƒ ¢ ˆ ’1+2anà ƒ ƒ ¢ ˆ ’2.
1
Expert's answer
2021-10-13T17:26:43-0400

Let us find the characteristic roots of the recurrence relation "a_n=a_{n-1}+2a_{n-2}." The characteristic equation "k^2=k+2" of the recurrence relation is equivalent to "k^2-k-2=0," and hence to "(k+1)(k-2)=0." It follows that the characteristic roots of the recurrence relation are "k_1=-1" and "k_2=2."